# Finite-state transducers for substitution tilings

### v1

@@ Abstract
We present a suite of algorithmic techniques for handling substitution tilings by treating a tile’s hierarchy of supertiles in a purely combinatorial fashion using finite state automata. The resulting techniques are very convenient for practical generation of patches of tilings such as hats, Spectres and Penrose tiles, both random and deliberately selected. They also permit some analyses of the represented tiling. A particular product of this process is two substitution systems for the hat tiling which are ‘unambiguous’ in that a single tile address uniquely determines the rest of the plane.
@@ 1 Introduction
Well-known aperiodic tilings such as the Penrose P2 and P3 tilings [Penrose1979], the Ammann-Beenker tiling ([10.5555/19304], Figures 10.4.14 and 10.4.15), the hat tiling [Smith_Jul2024] and the Spectre tiling [Smith_Sep2024], are generated by substitution systems, in which one tiling of the plane is converted (‘deflated’) into another, by consistently substituting a set of smaller tiles for each original tile. In some substitution systems, every level of this system looks the same; in others, a set of metatiles is repeatedly deflated, and then a different final substitution step turns the metatiles into the tiles of the desired end-product tiling.
@@
The substitution system itself provides an alternative way to identify each tile in a tiling of the plane, via an infinitely long address [Goodman-Strauss1999] describing the relationship of a particular tile $t_{0}$ to its supertile $t_{1}$ (both what type of tile $t_{1}$ is, and which subtile of $t_{1}$ is $t_{0}$), and the relationship of that in turn to the second-order supertile $t_{2}$, and so on. These addresses can be regarded as strings of symbols from a finite alphabet, and [Goodman-Strauss1999] suggests that finite state automata can be used to manipulate those strings. In particular an automaton is suggested which accepts the addresses of two tiles, one symbol at a time from each (as if they were combined by the zip operation in computer languages such as Python), and determines whether the two addresses describe tiles that are adjacent in the tiling.
This article develops that idea into a full system of algorithms for working with tilings in this way. The automaton described in [Goodman-Strauss1999] is modified slightly so that its input includes a choice of edge of each of the two input tiles, and reports not only whether the tiles could be adjacent at all, but whether they are adjacent along that specific pair of edges. A second automaton is constructed from this, which receives only a single edge address as input, and generates the address of its neighbour as output. This second automaton cannot always be constructed, but if it cannot, it is often (perhaps always) possible to modify the substitution system to remove the obstacle, with the modified system always being a ‘refinement’ of the old one – identical except that some tile types have been duplicated into multiple ‘subtypes’ carrying additional information about their role in the tiling. Algorithms for all of these constructions are presented.
@@
2 Combinatorial description of a substitution system
@@ 2.1 Sub-tiles, sub-edges, and adjacency map
As an initial example, we shall use a substitution system for the Penrose P2 tiling, in which the usual ‘kite’ and ‘dart’ tiles are each subdivided into two isosceles Robinson triangles [Robinson]. During deflation, each triangle is subdivided into either 2 or 3 smaller triangles, according to the rules shown in figure 1.
@@
During deflation, each supertile generates a number of subtiles; again we must assign each subtile an arbitrary index, and specify which tile type each subtile is. Indices can be anything you like as long as they are distinguishable; in this particular system, no two subtiles of a supertile have the same type, so we can use the subtile types themselves as indices. However, a computer implementation may prefer to keep things simple by requiring every tiling specification to index subtiles by consecutive integers.
@@
Finally, we must describe how adjacency of supertiles is related to adjacency of subtiles. If that $A$ supertile is deflated twice, so that its $B$ and $U$ subtiles each turn into two or three sub-subtiles, then the adjacency map for the $B$ subtile shows how its three sub-subtiles are arranged relative to each other, and similarly for the $U$, but to lay out the full set of sub-subtiles of the original $A$, we must also know which sub-tiles of the $B$ are adjacent to which of the $U$.
@@
Figure 2 shows an example: the full adjacency map for the deflation of a triangle of type $A$, and the geometric diagram it is derived from. The other adjacency maps are calculated similarly.
@@ 2.2 Spurs
We represent this in the combinatorial description of the tiling by allowing the adjacency map to contain entries which pair up two exterior edges (symbols of type $\mathbf{E}_{u,v}$), along with the entries already seen which pair up two interior edges, or one interior to one exterior edge. For example, in this diagram, the ‘upper’ side of the spur is the first of three sub-edges of the dart’s edge 1. Since those three sub-edges are indexed symmetrically as 2,0,+2-2,0,+2, the spur sub-edge will be $\mathbf{E}_{1,-2}$ in our notation. The ‘lower’ side of the spur is the single sub-edge of the dart’s edge 0. So the adjacency map for the dart will contain an entry pairing up $\mathbf{E}_{1,-2}\leftrightarrow\mathbf{E}_{0,0}$.
@@ 2.3 Edges are not always polygon edges
If ‘edge’ is taken to mean an edge of a polygon in the usual sense – a maximal straight segment of the shape’s boundary – then this property is not always satisfied. The hat monotile [Smith_Jul2024], regarded as a polygon, has 13 edges, consisting of six of length $1$, six of length $\sqrt{3}$, and one of length $2$. When it tiles the plane, the length-$2$ edge of the hat sometimes meets the length-$2$ edge of a neighbouring hat, but more often meets length-$1$ edges of two different hats, so that the midpoint of the length-$2$ edge is a point where three tiles meet. Figure 4 shows both of these possibilities.
@@
In general, for the purposes of a combinatorial description of this kind, one chooses how to divide the boundary of each tile into ‘edges’ by placing a vertex at any point on the boundary where it is possible for two other tiles to meet, and then propagating those vertices transitively to other tiles by placing a vertex anywhere that a neighbouring tile’s vertex can appear. Once this procedure has iterated to closure, the edges can be taken to be maximal segments of each tile’s boundary not containing a vertex.
@@ 2.4 Layers
Some substitution tilings use an auxiliary system of ‘metatiles’, which do not appear in the final tiling of the plane. One way to generate the hat tiling [Smith_Jul2024] uses a system of metatiles labelled $H,T,P,F$, with a set of substitution rules to deflate metatiles into other metatiles, and another set to deflate metatiles into hats. So one can generate a large patch of hat tiles by starting with a single metatile, say $H$; deflating it to a larger and larger patch of smaller sub-metatiles, for as many iterations as desired; and finally applying the other set of substitution rules just once, to turn each of the smallest $sub-\cdots$-sub-metatiles into a small number of actual hats.
@@
However, the period can be greater than 1: the layer structure of a substitution system need not terminate in a layer which is its own parent. An example is that the Robinson-triangle subdivision of the Penrose P2 tiling (kites and darts) can be deflated to the similar subdivision of the P3 tiling (thin and thick rhombs), and vice versa, creating a two-layer substitution system in which each layer’s parent is the other layer, and each deflation swaps between a P2 and P3 tiling of the plane. Some details of this are given in [DEBRUIJN198139].
@@ 2.5 Complete description of a combinatorial system
    • Number of sub-edges. For each edge of $t$, state how many sub-edges it turns into during deflation $P\to L$.
@@ 2.6 Geometric ‘source code’
All the algorithms described in this article can work using only the combinatorial representation of a tiling described above. However, that representation is not convenient for data entry. For more complicated substitution systems, when tiles have many edges and deflation of a tile creates many sub-tiles, the adjacency maps can be large and unwieldy (see Figures 45 and 47 for examples!). Also, since so much information about tile shape and position is missing from the combinatorial data, there is very little chance to cross-check it automatically to catch human error.
@@
A full example of this process is shown in a later section. Figure 49 shows the six tiles involved in the HTPF system for the hat tiling (the four metatiles, and the two handednesses of hat), with the metatiles using five distinct edge types and the hats two. Figures 50 and 51 show the edge deflation rules for the processes of deflating the metatiles into more metatiles, or to hats, respectively. Finally, Figures 52 and 53 show the result of deflating each tile’s outline according to each set of those rules, and the correct tiling of the interior of the patch with metatiles or hats respectively. (The edge types and directions are not shown, but if they were, the edge types in the patch outlines would correctly match those of the subtiles filling them.) Several of these diagrams show zero-thickness spurs, such as the one in the top left of the $H\to H$ deflation which matches $\mathbf{E}_{2,4}\leftrightarrow\mathbf{E}_{3,-2}$.
This geometric representation is more convenient for data entry than specifying the full combinatorial adjacency data directly. The user need only specify one entry of the adjacency map per sub-tile, and all the others are calculated automatically, by finding pairs of edges sharing both endpoints [^3] It is possible to find that a deflation diagram includes a pair of points $u,v$ such that there are multiple edges $u\to v$, making the adjacency map non-unique. Examples of this are shown in Figures 42 and 47. In this situation one expects the same number of edges from $u\to v$ as there are from $v\to u$, and then pairs them up arbitrarily. The different choices will only affect questions that do not change the output tiles, such as which side of one zero-thickness spur another one is considered to lie., and matching up each edge from $u\to v$ with one from $v\to u$.
@@ 2.7 Addresses of tiles and edges
In a plane tiling generated by a substitution system, we can assign to any tile an address [Goodman-Strauss1999] which gives its position within the combinatorial tile hierarchy.
@@ 3 Relation to regular languages
It is easy to see that $\mathcal{E}$ is a regular language, because a DFA can be constructed for it trivially, consisting of a special start state and a state for every $(\textit{layer},\textit{tile type})$ pair, with a transition between two states of the latter type whenever the supertile type and subtile index in a symbol are consistent with the starting state.
@@
Let $\mathcal{T}\times\mathcal{T}$ denote the set of ordered pairs of two tile addresses of the same length, represented as a sequence of ordered pairs containing one symbol from each address, as if they had been combined by Python zip. That is, if $(\sigma_{0},\sigma_{1},\ldots,\sigma_{n}),(\sigma^{\prime}_{0},\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n})\in\mathcal{T}$, then $((\sigma_{0},\sigma^{\prime}_{0}),(\sigma_{1},\sigma^{\prime}_{1}),\ldots,(\sigma_{n},\sigma^{\prime}_{n}))\in\mathcal{T}\times\mathcal{T}$.
Let $\mathcal{D}\subset\mathcal{T}\times\mathcal{T}$ be the ‘diagonal’ of $\mathcal{T}\times\mathcal{T}$: the language of strings made by zipping the same tile address with itself. This is trivially regular since $\mathcal{T}$ is.
@@
For every substitution system the author has so far investigated, $\mathcal{N}$ is a regular language. We do not have a proof of this for all substitution systems, but we can prove it subject to a plausible conjecture.
We begin by showing it for a restricted class of substitution systems: those without any zero-thickness spurs, such as the Robinson-triangle system for the P2 tiling shown in Figure 1.
@@
In a substitution system without spurs, $\mathcal{N}$ is a regular language.
@@
For each layer $L$ of the given substitution system, we can make a modified system which is identical except that its base layer is $L$, and define all the same languages described above for the modified system, which we shall denote $\mathcal{E}_{L},\mathcal{T}_{L},\mathcal{D}_{L},\mathcal{N}_{L}$. By the arguments above, $\mathcal{E}_{L},\mathcal{T}_{L},\mathcal{D}_{L}$ are regular for all $L$. We shall show that the same is true for $\mathcal{N}_{L}$.
Let $S\in\mathcal{N}_{L}$ begin with the symbols $\Bigl(\bigl(t_{0},e\bigr),\bigl(t^{\prime}_{0},e^{\prime}\bigr)\Bigr),\Bigl(\bigl(\xleftarrow{i_{1}}t_{1}\bigr),\bigl(\xleftarrow{i^{\prime}_{1}}t^{\prime}_{1}\bigr)\Bigr)$. That is, it describes two tiles of type $t_{0},t^{\prime}_{0}$, adjacent along edges $e,e^{\prime}$, with $t_{0}$ being the $i_{1}$th subtile of a supertile of type $t_{1}$, and similarly for $t^{\prime}_{0}$. Denote the ‘tail’ of $S$ after those first two symbols as $T$.
If $t_{1},t^{\prime}_{1}$ describe the same tile in the tiling, then its $i_{1}$th and $i^{\prime}_{1}$th subtiles must meet within that tile’s deflation diagram. In other words, in the adjacency map for tile type $t_{1}=t^{\prime}_{1}$, the entries $\mathbf{I}_{i_{1},e},\mathbf{I}_{i^{\prime}_{1},e^{\prime}}$ are mapped to each other. This is a valid string of $\mathcal{N}_{L}$ iff either $T=\epsilon$, or $(t_{1})\circ T\in\mathcal{D}_{P}$, where $P$ is the parent layer of $L$. (If any further symbols appear in $S$ after the common supertile $t_{1}$, then they must consist of a shared sequence of supertiles of $t_{1}$.)
Otherwise, the supertiles $t_{1},t^{\prime}_{1}$ of $t_{0},t^{\prime}_{0}$ are distinct. Since the substitution system has no spurs, the adjacency map for tile type $t_{1}$ must map $\mathbf{I}_{i_{1},e}$ to some $\mathbf{E}_{u,v}$, and the one for $t^{\prime}_{1}$ maps $\mathbf{I}_{i^{\prime}_{1},e^{\prime}}$ to $\mathbf{E}_{u^{\prime},v^{\prime}}$. For this to be a valid string of $\mathcal{N}_{L}$, we must have $v^{\prime}=-v$ so that the two subtiles appear in the same place along the edge between their supertiles, and the tail $T$ of the string must describe those supertiles being adjacent in turn, i.e. we need $\Bigl(\bigl(t_{1},u\bigr),\bigl(t^{\prime}_{1},u^{\prime}\bigr)\Bigr)\circ T\in\mathcal{N}_{P}$ for some $(t_{1},u,t^{\prime}_{1},u^{\prime},P)$.
It follows that the initial segment of $S$ until two supertiles coincide can be recognised by a DFA whose states correspond to the finite number of possible tuples $(t_{1},u,t^{\prime}_{1},u^{\prime},P)$, because if two strings $S,S^{\prime}\in\mathcal{N}_{L}$ can each be divided into head and tail segments $HT,H^{\prime},T^{\prime}$ passing through the same one of those states at the point of division, then $HT\in\mathcal{N}_{L}\Rightarrow\Bigl(\bigl(t_{1},u\bigr),\bigl(t^{\prime}_{1},u^{\prime}\bigr)\Bigr)\circ T\in\mathcal{N}_{P}\Rightarrow H^{\prime}T\in\mathcal{N}_{L}$, and similarly for $T^{\prime}$. So $\mathcal{N}_{L}$ is the concatenation of a string of that regular language, and one described by the regular language $\mathcal{D}_{P}$. ∎
The thin and thick rhombs from the Penrose P3 tiling, each shown with a deflation diagram turning them into two or three smaller rhombs. Both deflation diagrams include a zero-thickness spur. Since the rhombs are drawn as symmetric polygons but their behaviour is asymmetric, every rhomb is marked with an arrow indicating its orientation.

### v2

@@ Abstract
We present a suite of algorithmic techniques for handling substitution tilings by treating a tile’s hierarchy of supertiles in a purely combinatorial fashion using finite state automata. The resulting techniques are very convenient for practical generation of patches of aperiodic tilings such as hats, Spectres and Penrose tiles, both random and deliberately selected. They also permit some analyses of the represented tiling. A particular product of this process is two substitution systems for the hat tiling which are ‘unambiguous’ in that a single tile address uniquely determines the rest of the plane.
@@ 1 Introduction
Well-known aperiodic tilings such as the Penrose P2 and P3 tilings [10], the Ammann-Beenker tiling ([8], Figures 10.4.14 and 10.4.15), the hat tiling [12] and the Spectre tiling [11], are generated by substitution systems, in which one tiling of the plane is converted (‘deflated’) into another, by consistently substituting a set of smaller tiles for each original tile. In some substitution systems, every level of this system looks the same; in others, a set of metatiles is repeatedly deflated, and then a different final substitution step turns the metatiles into the tiles of the desired end-product tiling.
@@
The substitution system itself provides an alternative way to identify each tile in a tiling of the plane, via an infinitely long address [7] describing the relationship of a particular tile $t_{0}$ to its supertile $t_{1}$ (both what type of tile $t_{1}$ is, and which subtile of $t_{1}$ is $t_{0}$), and the relationship of that in turn to the second-order supertile $t_{2}$, and so on. These addresses can be regarded as strings of symbols from a finite alphabet, and [7] suggests that finite state automata can be used to manipulate those strings. In particular an automaton is suggested which accepts the addresses of two tiles, one symbol at a time from each (as if they were combined by the zip operation in programming languages such as Python), and determines whether the two addresses describe tiles that are adjacent in the tiling.
This article develops that idea into a full system of algorithms for working with tilings in this way. The automaton described in [7] is modified slightly so that its input includes a choice of edge of each of the two input tiles, and reports not only whether the tiles could be adjacent at all, but whether they are adjacent along that specific pair of edges. A second automaton is constructed from this, which receives only a single edge address as input, and generates the address of its neighbour as output. This second automaton cannot always be constructed, but if it cannot, the substitution system can be modified to remove the obstacle, with the modified system always being a ‘refinement’ of the old one – identical except that some tile types have been duplicated into multiple ‘subtypes’ carrying additional information about their role in the tiling. Algorithms for all of these constructions are presented.
@@
2 Specifying a substitution system
@@
2.1 Constraints
The term ‘substitution system’, or ‘substitution tiling’, has both general and specific usages.
In a general sense, a ‘substitution system’ might be any set of rules that transforms one family of tiles (or perhaps other objects) systematically into another. An example in this general sense might be a dissection of a single polygon into multiple polygons similar to itself, but of different sizes. Iterating this substitution leads to the number of differently sized polygons growing without bound. One might also choose to iterate the substitution a different number of times for each sub-polygon, e.g. continuing to substitute polygons as long as they exceed some threshold size.
Here, we are concerned with the more specific usage of ‘substitution system’ that applies to aperiodic tilings. For the purposes of this article, we expect a substitution system to meet some more stringent constraints:
Every tile is substituted the same number of times.
We only consider systems in which an entire tiling of the plane is transformed into an entire new tiling, by applying the substitution rules to every tile exactly once. That transformation can be iterated, so that each tile goes through substitution $n$ times, but we still expect that every tile has been substituted the same number of times.
Finitely many tile types up to congruence.
We consider systems with a finite set of ‘prototiles’ (reference subsets of the plane) fixed in advance, so that no matter how many times the substitution rules are applied, every tile remains congruent (not merely similar) to one of those prototiles. The example above violates this assumption because of the unbounded set of different sizes of tile.
Deflation makes more tiles.
We expect that after applying the substitution system to a whole tiling to create a deflated tiling, the total number of tiles is increased. This need not be true for a single tile in a single deflation, but it must be true that as $n\to\infty$, the total number of tiles in the $n$-times deflation of any starting tile increases without bound.
For example, there exists a substitution system that interchanges between the Penrose P2 and P3 tilings in each deflation, with the property that each P2 kite turns into just one P3 thick rhomb. But in the next deflation that P3 rhomb turns into two P2 tiles (a kite and a dart), and that number grows further as you iterate more times.
@@ 2.2 Sub-tiles, sub-edges, and adjacency map
As an initial example, we shall use a substitution system for the Penrose P2 tiling, in which the usual ‘kite’ and ‘dart’ tiles are each subdivided into two isosceles Robinson triangles [4]. During deflation, each triangle is subdivided into either 2 or 3 smaller triangles, according to the rules shown in figure 1.
@@
During deflation, each supertile generates a number of subtiles; again we must assign each subtile an arbitrary index, and specify which tile type each subtile is. Indices can be anything you like as long as they are distinguishable. Here we shall keep things simple for a computer implementation, by using integers $0,1,2,\ldots$ for the subtiles of each tile. [^1] This Robinson triangle system, and the corresponding one for the Penrose P3 tiling, both have the convenient property that no two subtiles of the same tile have the same type. As a result, you don’t really need a subtile index at all: you could reuse the subtile’s type as its index. But this property is rare and we shall not depend on it.
@@
Finally, we must describe how adjacency of supertiles is related to adjacency of subtiles. If that $A$ supertile is deflated twice, so that its $B$ and $U$ subtiles each turn into two or three sub-subtiles, then the adjacency map for the $B$ subtile shows how its three sub-subtiles are arranged relative to each other, and similarly for the $U$, but to lay out the full set of sub-subtiles of the original $A$, we must also know which sub-subtiles of the $B$ are adjacent to which of the $U$.
@@
Figure 2 shows an example: the full adjacency map for the deflation of a triangle of type $A$, and the geometric diagram it is derived from. The other adjacency maps are calculated similarly. The full set can be found in Figure 22.
@@ 2.3 Spurs
We represent this in the combinatorial description of the tiling by allowing the adjacency map to contain entries which pair up two exterior edges (symbols of type $\mathbf{E}_{u,v}$), along with the entries already seen which pair up two interior edges, or one interior to one exterior edge. For example, in this diagram, the ‘upper’ side of the spur is the first of three sub-edges of the dart’s edge 1. Since those three sub-edges are indexed symmetrically as 2,0,+2-2,0,+2, the spur sub-edge will be $\mathbf{E}_{1,-2}$ in our notation. The ‘lower’ side of the spur is the single sub-edge of the dart’s edge 0. So the adjacency map for the dart will contain an entry pairing up $\mathbf{E}_{1,-2}\leftrightarrow\mathbf{E}_{0,0}$. The full adjacency maps for this system, including this pairing of two external edges, can be seen in Figure 28.
@@ 2.4 Edges are not always polygon edges
If ‘edge’ is taken to mean an edge of a polygon in the usual sense – a maximal straight segment of the shape’s boundary – then this property is not always satisfied. The hat monotile [12], regarded as a polygon, has 13 edges, consisting of six of length $1$, six of length $\sqrt{3}$, and one of length $2$. When it tiles the plane, the length-$2$ edge of the hat sometimes meets the length-$2$ edge of a neighbouring hat, but more often meets length-$1$ edges of two different hats, so that the midpoint of the length-$2$ edge is a point where three tiles meet. Figure 4 shows both of these possibilities.
@@
In general, for the purposes of a combinatorial description of this kind, one chooses how to divide the boundary of each tile into ‘edges’ by placing a vertex at any point on the boundary where it is possible for two other tiles to meet, and then propagating those vertices transitively to other tiles by placing a vertex anywhere that a neighbouring tile’s vertex can appear. Once this procedure has iterated to closure, the edges can be taken to be maximal segments of each tile’s boundary not containing a vertex. (Additional vertices can be placed if desired for any other reason.)
@@ 2.5 Layers
Some substitution tilings use an auxiliary system of ‘metatiles’, which do not appear in the final tiling of the plane. One way to generate the hat tiling [12] uses a system of metatiles labelled $H,T,P,F$, with a set of substitution rules to deflate metatiles into other metatiles, and another set to deflate metatiles into hats. So one can generate a large patch of hat tiles by starting with a single metatile, say $H$; deflating it to a larger and larger patch of smaller sub-metatiles, for as many iterations as desired; and finally applying the other set of substitution rules just once, to turn each of the smallest $sub-\cdots$-sub-metatiles into a small number of actual hats.
@@
However, the period can be greater than 1: the layer structure of a substitution system need not terminate in a layer which is its own parent. An example, already mentioned in section 2.1, is that the Penrose P2 tiling (kites and darts) can be deflated to the P3 tiling (thin and thick rhombs), and vice versa, creating a two-layer substitution system in which each layer’s parent is the other layer, and each deflation swaps between a P2 and P3 tiling of the plane. (Some details of this are given in [5].)
@@ 2.6 Complete description of a combinatorial system
    • Number of sub-edges. For each edge of $t$, state how many sub-edges it turns into during deflation from $P\to L$.
@@ 2.7 Geometric ‘source code’
All the algorithms described in this article can work using only the combinatorial representation of a tiling described above. However, that representation is not convenient for data entry. For more complicated substitution systems, when tiles have many edges and deflation of a tile creates many sub-tiles, the adjacency maps can be large and unwieldy (see Figures 46 and 48 for examples!). Also, since so much information about tile shape and position is missing from the combinatorial data, there is very little chance to cross-check it automatically to catch human error.
@@
A full example of this process is shown in a later section. Figure 50 shows the six tiles involved in the HTPF system for the hat tiling (the four metatiles, and the two handednesses of hat), with the metatiles using five distinct edge types and the hats two. Figures 51 and 52 show the edge deflation rules for the processes of deflating the metatiles into more metatiles, or to hats, respectively. Finally, Figures 53 and 54 show the result of deflating each tile’s outline according to each set of those rules, and the correct tiling of the interior of the patch with metatiles or hats respectively. (The edge types and directions are not shown, but if they were, the edge types in the patch outlines would correctly match those of the subtiles filling them.) Several of these diagrams show zero-thickness spurs, such as the one in the top left of the $H\to H$ deflation which matches $\mathbf{E}_{2,4}\leftrightarrow\mathbf{E}_{3,-2}$.
This geometric representation is more convenient for data entry than specifying the full combinatorial adjacency data directly. The user need only specify one entry of the adjacency map per sub-tile, and all the others are calculated automatically, by finding pairs of edges sharing both endpoints [^4] It is possible to find that a deflation diagram includes a pair of points $u,v$ such that there are multiple edges $u\to v$, making the adjacency map non-unique. Examples of this are shown in Figures 43 and 48. In this situation one expects the same number of edges from $u\to v$ as there are from $v\to u$, and then pairs them up arbitrarily. The different choices will only affect questions that do not change the output tiles, such as which side of one zero-thickness spur another one is considered to lie., and matching up each edge from $u\to v$ with one from $v\to u$.
@@ 2.8 Addresses of tiles and edges
In a plane tiling generated by a substitution system, we can assign to any tile an address [7] which gives its position within the combinatorial tile hierarchy.
@@ 3 Relation to regular languages
It is easy to see that $\mathcal{E}$ is a regular language, because a DFA can be constructed for it trivially, consisting of a special start state and a state for every $(\textit{layer},\textit{tile type})$ pair, with a transition between two states of the latter type whenever the supertile type and subtile index in a symbol are consistent with the source state of the transition.
@@
Let $\mathcal{T}\times\mathcal{T}$ denote the set of ordered pairs of two tile addresses of the same length, represented as a sequence of ordered pairs containing one symbol from each address, as if they had been combined by Python zip. That is, if $(\sigma_{0},\sigma_{1},\ldots,\sigma_{n}),(\sigma^{\prime}_{0},\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n})\in\mathcal{T}$, then $((\sigma_{0},\sigma^{\prime}_{0}),(\sigma_{1},\sigma^{\prime}_{1}),\ldots,(\sigma_{n},\sigma^{\prime}_{n}))\in\mathcal{T}\times\mathcal{T}$. We define $\mathcal{E}\times\mathcal{E}$ similarly.
@@
$\mathcal{N}$ is a regular language.
@@
We begin by showing this for a restricted class of substitution systems: those in which the deflation of each tile type can be scaled to occupy the same region of the plane as the original tile, with no zero-thickness spurs or geometric distortion. An example is the Robinson-triangle system for the P2 tiling, shown in Figure 1.
If a substitution system has this property under a single deflation, then it has it under any finite number of deflations. So if two tiles $t,u$ are adjacent within the deflation of an order-$n$ supertile, then for every $0<i<n$, either their order-$i$ supertiles coincide, or are themselves adjacent along a whole edge, with the edge separating $t,u$ being part of that longer supertile edge.
Therefore, as we ascend from the tiles $(t,u)$ to their supertiles $(t_{1},u_{1}),\ldots,(t_{n},u_{n})$, each pair of supertiles is adjacent, until the first pair that coincide. After that, all further supertile pairs coincide.
This allows us to directly construct a state machine which matches $\mathcal{N}$. For the initial segment of a string before two supertiles coincide, each state $S(L,t,u,e,f)$ describes a layer $L$ of the system; a pair of tile types $t,u$ within that layer; and a pair of edge indices $e,f$ of those tiles, indicating their adjacency: edge $e$ of tile $t$ meets edge $f$ of tile $u$.
Transitions can be determined using the adjacency maps, such as the one shown in Figure 2. In a state $S(L,t,u,e,f)$, if there exist tile types $t^{\prime},u^{\prime}$ in $L$’s parent layer $P$, and indices $i,j,k,e^{\prime},f^{\prime}$ such that
  • subtile $i$ of tile type $t^{\prime}$ (when deflated from $P\to L$) is type $t$
  • subtile $j$ of tile type $u^{\prime}$, similarly, is type $u$
  • the adjacency map for deflating tile $t^{\prime}$ maps $\mathbf{I}_{i,e}\leftrightarrow\mathbf{E}_{e^{\prime},k}$
  • the adjacency map for deflating tile $u^{\prime}$ maps $\mathbf{I}_{j,f}\leftrightarrow\mathbf{E}_{f^{\prime},-k}$
then this is saying that $t,u$ occupy compatible positions along the boundaries of two edges of $t^{\prime},u^{\prime}$, such that if edge $e^{\prime}$ of $t^{\prime}$ meets edge $f^{\prime}$ of $u^{\prime}$, then their subtiles would meet in the way described by the original state $S(L,t,u,e,f)$. Hence, a transition exists from $S(L,t,u,e,f)\to S(P,t^{\prime},u^{\prime},e^{\prime},f^{\prime})$, on a symbol pair $\Bigl(\bigl(\xleftarrow{i}t^{\prime}\bigr),\bigl(\xleftarrow{j}u^{\prime}\bigr)\Bigr)$.
At the point where the supertiles of $t,u$ coincide, we enter a second family of states, describing a single tile in a particular layer, say $A(L,t)$. These are the accepting states of our automaton. The first transition into an accepting state occurs when subtiles $i,j$ of a supertile type $s$ match the types $t,u$ from an existing state, and those subtiles meet internally to the supertile’s deflation, along the correct edges. That is, if $s$’s adjacency map includes the entry $\mathbf{I}_{i,e}\leftrightarrow\mathbf{I}_{j,f}$, then a transition exists from $S(L,t,u,e,f)\to A(P,s)$ on the symbol pair $\Bigl(\bigl(\xleftarrow{i}s\bigr),\bigl(\xleftarrow{j}s\bigr)\Bigr)$. Thereafter, transitions between accepting states require both symbols in the pair to be identical, and to continue to spell out a valid tile address, exactly as in the natural DFA matching $\mathcal{T}$.
Now we must extend this construction to more general substitution systems: those in which tile edges deflate to non-straight outlines, and especially those with zero-thickness spurs. In these systems, it need not be true that two tiles $t,u$ sharing an edge have supertiles $t^{\prime},u^{\prime}$ also sharing an edge.
Increasingly complex outlines derived from the P2 kite and dart shapes, in each step transforming the edges via the deflation map. After many iterations, each shape has converged to an irregular fractal.